# Prospects for Probing Sub-GeV Leptophilic Dark Matter with the Future VLAST

### v1

@@ IV Allowed parameter space
In Fig. 3, we show the $95\%$ allowed region projected onto the ($m_{\text{DM}}$, $\bar{\sigma}_{e}$) plane for electrophilic interactions. The light blue regions indicate the non-resonance regime, which includes the forbidden and $m_{\rm DM}>m_{\rm MED}$ regions (as shown in Fig. 2). Our scan for the mediator mass is above $1\,\rm{MeV}$, specifically $m_{\rm{MED}}\gg\alpha m_{e}$, allowing us to use the heavy mediator form-factor $F_{\rm{DM}}(q)=1$. It is worth noting that most of the non-resonance region in both the $p$-wave (middle panel) and $(s+p)$-wave (right panel) scenarios have already been excluded by XENON1T Aprile et al. (2019), while the resonance region exhibits a suppressed $\bar{\sigma}_{e}$, making it difficult to be probed with XENON1T. This suppression arises from the Planck relic density constraint, which imposes an upper limit on the coupling constants $g_{D}$ and $g_{f}$. In the following, we will focus exclusively on the resonance region, which can bypass the constraints imposed by relic density, CMB, and DM direct detection.
@@
Figure 4: The $95\%$ allowed regions for the resonance regions. The coupling product $M_{D}\times g_{e\,(\mu)}$ (left column) and $g_{D}\times g_{e\,(\mu)}$ (middle and right columns) are the function of the resonance parameter $\xi$ in the resonance regime. The resonance parameter is defined as $\xi=\sqrt{(m_{\rm{MED}}^{2}-4m_{\rm{DM}}^{2})/4m_{\rm{DM}}^{2}}$.
@@
  • The upper bound on $\xi$ is determined by the resonance condition. Ref. Wang et al. (2025) suggests that $\xi^{2}$ can be as small as $\mathcal{O}(10^{-16})$, but achieving such a small value would require extreme fine-tuning, which is beyond the scope of this work.
@@ V Probing resonance region by VLAST
Figure 5: Expected sensitivity of VLAST to DM annihilation via mediator resonance in the Draco dwarf galaxy over a 5-year observation period. The red and orange regions denote the parameter space accessible to VLAST for the electrophilic (upper panels) and muonphilic (lower panels) DM, respectively, while the gray region represents the parameter space that remains undetectable.
In Fig. 5, we show the projected sensitivity of VLAST to DM annihilation via mediator resonance within the Draco dwarf galaxy over a 5-year observation period, presenting three different interactions across three columns for electrophilic (upper panels) and muonphilic (lower panels) DM. The red and orange regions indicate the parameter space where VLAST is sensitive to DM annihilation signals, while the gray region represents the parameter space where the annihilation cross-section is too low for the VLAST future detection.
For annihilation via mediator resonance in the $p$-wave and $(s+p)$-wave cases, VLAST is expected to provide strong sensitivity to DM annihilation signals. However, VLAST exhibits limited sensitivity to the $s$-wave scenario, even with the resonance mechanism. This limitation arises from the combination of the DM relic abundance and CMB constraints: achieving the correct relic density requires a sufficiently large annihilation cross-section, while evading stringent CMB constraints needs a small cross-section during the CMB epoch. The only way to meet both conditions in the $s$-wave case is for the resonance peak to be near the freeze-out epoch, which generates a suppression of the annihilation cross-section at later times when the DM velocity is sufficiently lower. In the Draco dwarf galaxy, where the typical DM relative velocity is approximately $\sim 10^{-4}$ in the unit of light speed, this suppression persists. Therefore, DM annihilation with $s$-wave resonance remains undetectable.
@@ VI Summary and conclusion
A central focus of our study is the resonance region (where $m_{\text{MED}}\approx 2m_{\text{DM}}$), which allows for enhanced annihilation cross-sections during DM freeze-out without violating CMB constraints in the present epoch. This resonance condition enables viable $s$-wave, $p$-wave, and mixed $(s+p)$-wave annihilation, although the parameter space remains heavily constrained for the $s$-wave case due to its velocity-independent behavior.
@@
We further distinguish between electrophilic and muonphilic interactions. In the electrophilic case, direct detection experiments impose strong constraints on the non-resonance parameter space due to tree-level DM-electron scattering. However, VLAST can effectively probe the resonance region that escapes these bounds. The muonphilic case, in contrast, avoids strong direct detection limits due to suppressed loop-level interactions with electrons, making the muon channel an attractive target for future indirect searches.
@@ Acknowledgments
We were supported by the National Key Research and Development Program of China (No. 2022YFF0503304), the Project for Young Scientists in Basic Research of the Chinese Academy of Sciences (No. YSBR-092). TPT is also supported by the Jiangsu Province Post Doctoral Foundation (No. 2024ZB713).

### v2

@@ IV Allowed parameter space
In Fig. 3, we show the $2\sigma$ allowed region projected onto the ($m_{\text{DM}}$, $\bar{\sigma}_{e}$) plane for electrophilic interactions. The light blue regions indicate the non-resonance regime, which includes the forbidden and $m_{\rm DM}>m_{\rm MED}$ regions (as shown in Fig. 2). Our scan for the mediator mass is above $1\,\rm{MeV}$, specifically $m_{\rm{MED}}\gg\alpha m_{e}$, allowing us to use the heavy mediator form-factor $F_{\rm{DM}}(q)=1$. It is worth noting that most of the non-resonance region in both the $p$-wave (middle panel) and $(s+p)$-wave (right panel) scenarios have already been excluded by the null results from XENON1T XENON:2019gfn, DarkSide-50 DarkSide:2022knj, and DAMIC-M DAMIC-M:2023gxo, while the resonance region exhibits a suppressed $\bar{\sigma}_{e}$, making it difficult to be probed with these experiments. This suppression arises from the Planck relic density constraint, which imposes an upper limit on the coupling constants $g_{D}$ and $g_{f}$. In the following, we will focus exclusively on the resonance region, which can bypass the constraints imposed by relic density, CMB, and DM direct detection.
@@
Figure 4: The $2\sigma$ allowed regions for the resonance regions. The coupling product $M_{D}\times g_{e\,(\mu)}$ (left column) and $g_{D}\times g_{e\,(\mu)}$ (middle and right columns) are the function of the resonance parameter $\xi$ in the resonance regime. The resonance parameter is defined as $\xi=\sqrt{(m_{\rm{MED}}^{2}-4m_{\rm{DM}}^{2})/4m_{\rm{DM}}^{2}}$.
@@
  • The upper bound on $\xi$ is determined by the resonance condition. Ref. Wang:2025tdx suggests that $\xi^{2}$ can be as small as $\mathcal{O}(10^{-16})$, but achieving such a small value would require extreme fine-tuning, which is beyond the scope of this work.
@@ V Probing resonance region by VLAST
Figure 5: Expected sensitivity of VLAST to DM annihilation via mediator resonance in Draco dwarf galaxy over a 5-year observation period. The red and orange regions denote the parameter space accessible to VLAST for the electrophilic (upper panels) and muonphilic (lower panels) DM, respectively, while the gray region represents the parameter space that remains undetectable.
In Fig. 5, we show the projected sensitivity of VLAST to DM annihilation via mediator resonance within Draco dwarf galaxy over a 5-year observation period, presenting three different interactions across three columns for electrophilic (upper panels) and muonphilic (lower panels) DM. The red and orange regions indicate the parameter space where VLAST is sensitive to DM annihilation signals, while the gray region represents the parameter space where the annihilation cross section is too low for the VLAST future detection.
For annihilation via mediator resonance in the $p$-wave and $(s+p)$-wave cases, VLAST is expected to provide strong sensitivity to DM annihilation signals. The white regions in the electrophilic $p$-wave (upper middle panel) and $(s+p)$-wave (upper right panel) correspond to parameter points that are excluded by existing constraints rather than being undetectable. At very low masses ($m_{\rm DM}\lesssim 3\,\rm{MeV}$), the lower boundary is set by the $\Delta N_{\rm eff}$ constraint. For DM masses between roughly $3\,\rm{MeV}$ and $55\,\rm{MeV}$, parameter points with large annihilation cross sections are excluded by CMB constraints, as the Planck limits on energy injection become increasingly stringent toward lighter DM masses.
VLAST exhibits limited sensitivity to the $s$-wave scenario, even with the resonance mechanism. This limitation arises from the combination of the DM relic abundance and CMB constraints: achieving the correct relic density requires a sufficiently large annihilation cross section, while evading stringent CMB constraints needs a small cross section during the CMB epoch. The only way to meet both conditions in the $s$-wave case is for the resonance peak to be near the freeze-out epoch, which generates a suppression of the annihilation cross section at later times when the DM velocity is sufficiently lower. Consequently, the white regions in the $s$-wave panels (upper and lower left) correspond to parameter points excluded by this tension between relic density and CMB constraints. In Draco dwarf galaxy, where the typical DM relative velocity is approximately $\sim 10^{-4}$ in the unit of light speed, this suppression persists. Therefore, DM annihilation with $s$-wave resonance remains undetectable.
@@ VI Summary and conclusion
A central focus of our study is the resonance region (where $m_{\text{MED}}\approx 2m_{\text{DM}}$), which allows for enhanced annihilation cross sections during DM freeze-out without violating CMB constraints in the present epoch. This resonance condition enables viable $s$-wave, $p$-wave, and mixed $(s+p)$-wave annihilation, although the parameter space remains heavily constrained for the $s$-wave case due to its velocity-independent behavior.
@@
We further distinguish between electrophilic and muonphilic interactions. In the electrophilic case, direct detection experiments impose strong constraints on the non-resonance parameter space due to tree-level DM-electron scattering. However, VLAST can effectively probe the resonance region that escapes these bounds. In contrast, the muonophilic case avoids stringent direct detection limits by suppressing loop-level interactions with electrons, making the muon channel an appealing annihilation final state for future indirect searches.
@@ Acknowledgments
We were supported by the National Key Research and Development Program of China (No. 2022YFF0503304), and the Project for Young Scientists in Basic Research of the Chinese Academy of Sciences (No. YSBR-092). TPT is also supported by the Jiangsu Province Post Doctoral Foundation (No. 2024ZB713).